A Complete Bibliography of Internet STD Documents

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

19 July 2005 Version 1.18

Title word cross-reference

1 [1]. 10 [10]. 10646 [63]. 11 [11]. 12 [12]. 13 [13]. 14 [14]. 15 [15]. 16 [16]. 17 [17]. 18 [18]. 19 [19].

2 [2]. 20 [20]. 21 [21]. 22 [22]. 23 [23]. 24 [24]. 25 [25]. 26 [26]. 27 [27]. 28 [28]. 29 [29].

3 [3, 53]. 30 [30]. 31 [31]. 32 [32]. 33 [33]. 34 [34]. 35 [35]. 36 [36]. 37 [37]. 38 [38]. 39 [39].

4 [4]. 40 [40]. 41 [41]. 42 [42]. 43 [43]. 44 [44]. 45 [45]. 46 [46]. 47 [47]. 48 [48]. 48.bit [37]. 49 [49].

5 [5]. 51 [51]. 52 [52]. 53 [53]. 54 [54]. 55 [55]. 56 [56]. 57 [57]. 58 [58]. 59 [59].

6 [6]. 60 [60]. 61 [61]. 62 [62]. 63 [63]. 64 [64]. 65 [65]. 66 [66]. 7 [7]. 8 [8, 63]. 802 [43]. 9 [9]. Access [40]. Active [24]. Address [37, 38]. Addresses [37]. Ahead [29]. Applicability [57]. Applications [64]. ARCNET [46]. ARP [36]. ARPA [11]. Assigned [2]. Audio [65]. Base [17, 59]. based [16, 17]. Binary [27]. Character [22]. Command [60]. Communication [3]. Concepts [13, 19]. Conferences [65]. Control [7, 65]. Converting [37]. DARPA [5, 7]. Datagram [6]. Datagrams [41, 43, 48, 52, 42, 47]. Day [23]. Daytime [25]. DCN [44]. Definitions [50]. Discard [21]. Domain [13, 14]. Echo [28, 20]. Ethernet [37, 41, 42, 50]. Ethernet-like [50]. experimental [42]. Extended [32]. Extension [60]. Exterior [18]. Facilities [13]. FDDI [36]. File [9]. Formal [18]. format [11, 63]. Frame [55]. Gateway [18]. Gateways [4]. Generator [22]. Generic [66]. Go [29]. Hardware [37]. Host [39, 40]. Hosts [3]. HYPERchannel [45]. Identification [16]. Identifier [66]. IEEE [43]. II [17]. IMP [39]. Implementation [12]. Information [16, 17, 34, 59, 58]. Interconnect [55]. Interconnection [39]. Interface [50, 39]. Internet [3, 4, 7, 1, 5, 11, 45]. Internets [16, 17]. IP [36, 41, 42, 43, 46, 47, 48, 52]. IPX [49]. ISO [35, 63]. Layers [3]. like [50]. lines [47]. List [32]. Local [44]. Local-Network [44]. Mail [10, 14]. Managed [50]. Management [15, 16, 58, 59, 62, 17]. Mark [31]. Message [39]. messages [11]. Methods [19]. MIB [17]. MIB-II [17]. Minimal [65]. Monitoring [59]. Multiprotocol [55]. Names [13]. NetBIOS [48]. Network [15, 17, 37, 44, 59, 62, 12, 45]. Networks [36, 41, 42, 43, 48, 49, 46]. Nonstandard [47]. Numbers [2].

Objects [50]. Office [53]. Official [1]. One [61]. One-Time [61]. Option [28, 29, 30, 31, 32]. Options [32]. OSPF [54].

Packets [49]. Password [61]. IP-based [17, 16]. UDP [19]. Pipelining [60]. Point [51]. Point-to-Point [51]. Post [53]. PPP [51]. Processor [39]. Profile [65]. Program [5, 7]. Protocol [1, 5, 7, 10, 12, 18, 33, 37, 53, 57, 62, 64, 6, 8, 9, 15, 19, 20, 21, 22, 23, 25, 26, 34, 38, 40, 45, 51]. Protocols [44].

Quote [23].

Real [64]. Real-Time [64]. Relay [55]. Remote [59]. Requirements [3, 4]. Reserved [50]. Resolution [37, 38]. Resource [66]. Reverse [38]. Revised [39]. Revision [33]. RIP [56, 57]. Routing [14, 34]. RTP [64, 65].

serial [47]. Service [19, 52, 60, 35]. Simple [10, 15, 62]. SLIP [47]. SMDS [52]. SMIv2 [58]. SMTP [60]. SNMPv3 [62]. Specification [12, 45, 40, 5, 7, 18]. Specifications [39]. Standard [11, 19, 41, 42, 43, 48, 49]. Standards [1]. Statement [57]. Status [30]. STD [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. Structure [16, 58]. Suppress [29]. Syntax [66]. System [14, 45, 61].

TCP [17, 19, 16, 35]. TCP/IP [17, 16]. TCP/IP-based [17, 16]. TCP/ UDP [19]. Telnet [8, 27, 28, 29, 30, 31, 32]. text [11]. TFTP [33]. Time [12, 61, 64, 26]. Timing [31]. top [35]. traffic [46]. Transfer [9, 10]. transformation [63]. Transmission [36, 37, 41, 43, 48, 49, 52, 42, 47, 7, 27]. Transmitting [46]. Transport [19, 35, 64]. Types [50].

Uniform [66]. URI [66]. User [6]. Users [24]. UTF [63]. UTF-8 [63].

Version [12, 35, 53, 54, 56, 57, 58, 62]. Video [65].

References

STD0001

[1] J. Reynolds and S. Ginoza. STD 1: Internet Official Protocol Standards,

    July 2004. URL ftp://ftp.isi.edu/in-notes/rfc3600.txt,   ftp:/
    /ftp.isi.edu/in-notes/rfc3700.txt,   ftp://ftp.math.utah.edu/
    pub/rfc/rfc3600.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc3700.

txt. See also RFC3700 [?]. Obsoletes RFC3600 [?].

STD0002

[2] J. Reynolds and J. Postel. STD 2: Assigned numbers, October 1994.

URL ftp://ftp.isi.edu/in-notes/rfc1340.txt;   ftp://ftp.isi.
edu/in-notes/rfc1700.txt;   ftp://ftp.isi.edu/in-notes/rfc3232.
txt;   ftp://ftp.isi.edu/in-notes/std/std2.txt;   ftp://ftp.math.
utah.edu/pub/rfc/rfc1340.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc1700.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc3232.txt;
ftp://ftp.math.utah.edu/pub/rfc/std/std2.txt. See also RFC1700

[?]. Obsoleted by RFC3232 [?]. Obsoletes RFC1340 [?].

STD0003

[3] R. Braden. STD 3: Requirements for internet hosts — communication lay-

ers, October 1989. URL ftp://ftp.isi.edu/in-notes/rfc1122.txt;
 ftp://ftp.isi.edu/in-notes/rfc1123.txt;   ftp://ftp.isi.edu/
in-notes/std/std3.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1122.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1123.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std3.txt. See also RFC1122, RFC1123

[?, ?].

STD0004

[4] R. Braden and J. Postel. STD 4: Requirements for internet gateways, June

1987. URL ftp://ftp.isi.edu/in-notes/rfc1009.txt;   ftp://ftp.
isi.edu/in-notes/std/std4.txt;   ftp://ftp.math.utah.edu/pub/
rfc/rfc1009.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std4.txt.

See also RFC1009 [?].

STD0005

[5] J. Postel. STD 5: Internet Protocol: DARPA Internet Program Protocol Specification, September 1981. URL ftp://ftp.isi.edu/in-notes/

rfc1112.txt;   ftp://ftp.isi.edu/in-notes/rfc791.txt; ftp://ftp.
isi.edu/in-notes/rfc792.txt;   ftp://ftp.isi.edu/in-notes/rfc919.
txt; ftp://ftp.isi.edu/in-notes/rfc922.txt;   ftp://ftp.isi.edu/
in-notes/rfc950.txt;   ftp://ftp.isi.edu/in-notes/std/std5.txt;
 ftp://ftp.math.utah.edu/pub/rfc/rfc1112.txt;   ftp://ftp.math.
utah.edu/pub/rfc/rfc791.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc792.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc919.txt;   ftp:
//ftp.math.utah.edu/pub/rfc/rfc922.txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc950.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/

std5.txt. See also RFC0791, RFC0792, RFC0919, RFC0922, RFC0950, RFC1112 [?, ?, ?, ?, ?, ?].

STD0006

[6] J. Postel. STD 6: User Datagram Protocol, August 1980. URL ftp://ftp.

isi.edu/in-notes/rfc768.txt;   ftp://ftp.isi.edu/in-notes/std/
std6.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc768.txt;   ftp:
//ftp.math.utah.edu/pub/rfc/std/std6.txt. See also RFC0768 [?].

STD0007

[7] J. Postel. STD 7: Transmission Control Protocol: DARPA Internet Program Protocol Specification, September 1981. URL ftp://ftp.isi.edu/

in-notes/rfc793.txt;   ftp://ftp.isi.edu/in-notes/std/std7.txt;
 ftp://ftp.math.utah.edu/pub/rfc/rfc793.txt;   ftp://ftp.math.
utah.edu/pub/rfc/std/std7.txt. See also RFC0793 [?].

STD0008

[8] J. Postel and J. Reynolds. STD 8: Telnet Protocol, May 1983.

URL ftp://ftp.isi.edu/in-notes/rfc854.txt;   ftp://ftp.isi.
edu/in-notes/rfc855.txt;   ftp://ftp.isi.edu/in-notes/std/std8.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc854.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc855.txt;   ftp://ftp.math.utah.edu/pub/

rfc/std/std8.txt. See also RFC0854, RFC0855 [?, ?].

STD0009

[9] J. Postel and J. Reynolds. STD 9: File Transfer Protocol, October 1985.

URL ftp://ftp.isi.edu/in-notes/rfc2228.txt;   ftp://ftp.isi.
edu/in-notes/rfc2640.txt;   ftp://ftp.isi.edu/in-notes/rfc959.
txt;   ftp://ftp.isi.edu/in-notes/std/std9.txt;   ftp://ftp.math.
utah.edu/pub/rfc/rfc2228.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc2640.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc959.txt;   ftp:
//ftp.math.utah.edu/pub/rfc/std/std9.txt. See also RFC0959 [?].

Updated by RFC2228, RFC2640 [?, ?].

STD0010

[10] Jonathan B. Postel. STD 10: Simple mail transfer protocol, August 1982.

URL ftp://ftp.isi.edu/in-notes/rfc1869.txt;   ftp://ftp.isi.
edu/in-notes/rfc2821.txt;   ftp://ftp.isi.edu/in-notes/rfc772.
txt; ftp://ftp.isi.edu/in-notes/rfc780.txt;   ftp://ftp.isi.edu/
in-notes/rfc788.txt; ftp://ftp.isi.edu/in-notes/rfc821.txt;
ftp://ftp.isi.edu/in-notes/rfc974.txt;   ftp://ftp.isi.edu/in-
notes/std/std10.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1869.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc2821.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc772.txt;   ftp://ftp.math.utah.edu/pub/
rfc/rfc780.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc788.txt;

 ftp://ftp.math.utah.edu/pub/rfc/rfc821.txt;   ftp://ftp.math.
utah.edu/pub/rfc/rfc974.txt;   ftp://ftp.math.utah.edu/pub/rfc/

std/std10.txt. See also RFC0821, RFC1869. RFC974 [?, ?, ?]. Obsoleted by RFC2821 [?]. Obsoletes RFC788, RFC780, RFC772 [?, ?, ?].

STD0011

[11] David H. Crocker. STD 11: Standard for the format of ARPA Internet text messages, August 13, 1982. URL ftp://ftp.isi.edu/in-notes/

rfc2822.txt;   ftp://ftp.isi.edu/in-notes/rfc733.txt; ftp://ftp.
isi.edu/in-notes/rfc822.txt;   ftp://ftp.isi.edu/in-notes/std/
std11.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc2822.txt;   ftp:
//ftp.math.utah.edu/pub/rfc/rfc733.txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc822.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/

std11.txt. See also RFC0822 [?]. Obsoleted by RFC2822 [?]. Obsoletes RFC0733 [?].

STD0012

[12] David L. Mills. STD 12: Network Time Protocol (Version 2) Specification and Implementation, September 1989. URL ftp://ftp.isi.edu/in-

notes/rfc1119.ps;   ftp://ftp.isi.edu/in-notes/rfc1119.txt;
ftp://ftp.isi.edu/in-notes/std/std12.txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc1119.ps;   ftp://ftp.math.utah.edu/pub/rfc/rfc1119.
txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std12.txt. See also

RFC1119 [?].

STD0013

[13] P. Mockapetris. STD 13: Domain Names — Concepts and Facili-

ties, November 1987. URL ftp://ftp.isi.edu/in-notes/rfc1034.txt;
 ftp://ftp.isi.edu/in-notes/rfc1035.txt;   ftp://ftp.isi.edu/
in-notes/std/std13.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1034.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1035.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std13.txt. See also RFC1034, RFC1035

[?, ?].

STD0014

[14] Craig Partridge. STD 14: Mail routing and the Domain sys-

tem, January 1986.  URL ftp://ftp.isi.edu/in-notes/rfc974.
txt;   ftp://ftp.isi.edu/in-notes/std/std14.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc974.txt;   ftp://ftp.math.utah.edu/pub/

rfc/std/std14.txt. See also RFC0974 [?].

STD0015

[15] J. Case, M. Fedor, M. Schoffstall, and J. Davin. STD 15: A Simple Network Management Protocol, May 1990. URL ftp://ftp.isi.edu/

in-notes/rfc1157.txt;   ftp://ftp.isi.edu/in-notes/std/std15.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1157.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std15.txt. See also RFC1157 [?].

STD0016

[16] M. Rose and K. McCloghrie. STD 16: Structure and identification of management information for TCP/IP-based internets, May 1990.

URL ftp://ftp.isi.edu/in-notes/rfc1065.txt;   ftp://ftp.isi.
edu/in-notes/rfc1155.txt;   ftp://ftp.isi.edu/in-notes/rfc1212.
txt;   ftp://ftp.isi.edu/in-notes/std/std16.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc1065.txt;   ftp://ftp.math.utah.edu/
pub/rfc/rfc1155.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1212.
txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std16.txt. See also

RFC1155, RFC1212 [?, ?]. Obsoletes RFC1065 [?].

STD0017

[17] K. McCloghrie and M. Rose. STD 17: Management Information Base for Network Management of TCP/IP-based internets: MIB-

II, March 1991. URL ftp://ftp.isi.edu/in-notes/rfc1158.txt;
 ftp://ftp.isi.edu/in-notes/rfc1213.txt;   ftp://ftp.isi.edu/
in-notes/std/std17.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1158.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1213.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std17.txt. See also RFC1213 [?]. Obso-

letes RFC1158 [?].

STD0018

[18] D. L. Mills. STD 18: Exterior Gateway Protocol Formal Spec-

ification, April 1984. URL ftp://ftp.isi.edu/in-notes/rfc904.
txt;   ftp://ftp.isi.edu/in-notes/std/std18.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc904.txt;   ftp://ftp.math.utah.edu/pub/

rfc/std/std18.txt. See also RFC0904 [?].

STD0019

[19] NetBIOS Working Group. STD 19: Protocol Standard for a NetBIOS Service on a TCP/UDP Transport: Concepts and Methods, March 1987.

URL ftp://ftp.isi.edu/in-notes/rfc1001.txt;   ftp://ftp.isi.
edu/in-notes/rfc1002.txt;   ftp://ftp.isi.edu/in-notes/std/std19.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1001.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc1002.txt;   ftp://ftp.math.utah.edu/

pub/rfc/std/std19.txt. See also RFC1001, RFC1002 [?, ?].

STD0020

[20] J. Postel. STD 20: Echo Protocol, May 1983. URL ftp://ftp.isi.edu/

in-notes/rfc862.txt;   ftp://ftp.isi.edu/in-notes/std/std20.txt;
 ftp://ftp.math.utah.edu/pub/rfc/rfc862.txt;   ftp://ftp.math.
utah.edu/pub/rfc/std/std20.txt. See also RFC0862 [?].

STD0021

[21] J. Postel. STD 21: Discard Protocol, May 1983. URL ftp://ftp.isi.edu/

in-notes/rfc863.txt;   ftp://ftp.isi.edu/in-notes/std/std21.txt;
 ftp://ftp.math.utah.edu/pub/rfc/rfc863.txt;   ftp://ftp.math.
utah.edu/pub/rfc/std/std21.txt. See also RFC0863 [?].

STD0022

[22] J. Postel. STD 22: Character Generator Protocol, May 1983.

URL ftp://ftp.isi.edu/in-notes/rfc864.txt;   ftp://ftp.isi.
edu/in-notes/std/std22.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc864.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std22.txt. See

also RFC0864 [?].

STD0023

[23] J. Postel. STD 23: Quote of the Day Protocol, May 1983. URL ftp://ftp.

isi.edu/in-notes/rfc865.txt;   ftp://ftp.isi.edu/in-notes/std/
std23.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc865.txt;   ftp:
//ftp.math.utah.edu/pub/rfc/std/std23.txt. See also RFC0865 [?].

STD0024

[24] J. Postel. STD 24: Active Users, May 1983. URL ftp://ftp.isi.edu/in-

notes/rfc866.txt;   ftp://ftp.isi.edu/in-notes/std/std24.txt;
 ftp://ftp.math.utah.edu/pub/rfc/rfc866.txt;   ftp://ftp.math.
utah.edu/pub/rfc/std/std24.txt. See also RFC0866 [?].

STD0025

[25] J. Postel. STD 25: Daytime Protocol, May 1983. URL ftp://ftp.isi.

edu/in-notes/rfc867.txt;   ftp://ftp.isi.edu/in-notes/std/std25.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc867.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std25.txt. See also RFC0867 [?].

STD0026

[26] J. Postel and K. Harrenstien. STD 26: Time Protocol, May 1983.

URL ftp://ftp.isi.edu/in-notes/rfc868.txt;   ftp://ftp.isi.
edu/in-notes/std/std26.txt;   ftp://ftp.math.utah.edu/pub/rfc/

rfc868.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std26.txt. See

also RFC0868 [?].

STD0027

[27] J. Postel and J. Reynolds. STD 27: Telnet Binary Transmission, May 1983.

URL ftp://ftp.isi.edu/in-notes/rfc856.txt;   ftp://ftp.isi.
edu/in-notes/std/std27.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc856.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std27.txt. See

also RFC0856 [?].

STD0028

[28] J. Postel and J. Reynolds. STD 28: Telnet Echo Option, May 1983.

URL ftp://ftp.isi.edu/in-notes/rfc857.txt;   ftp://ftp.isi.
edu/in-notes/std/std28.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc857.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std28.txt. See

also RFC0857 [?].

STD0029

[29] J. Postel and J. Reynolds. STD 29: Telnet Suppress Go Ahead

Option, May 1983.  URL ftp://ftp.isi.edu/in-notes/rfc858.
txt;   ftp://ftp.isi.edu/in-notes/std/std29.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc858.txt;   ftp://ftp.math.utah.edu/pub/

rfc/std/std29.txt. See also RFC0858 [?].

STD0030

[30] J. Postel and J. Reynolds. STD 30: Telnet Status Option, May 1983.

URL ftp://ftp.isi.edu/in-notes/rfc859.txt;   ftp://ftp.isi.
edu/in-notes/std/std30.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc859.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std30.txt. See

also RFC0859 [?].

STD0031

[31] J. Postel and J. Reynolds. STD 31: Telnet Timing Mark Option, May 1983.

URL ftp://ftp.isi.edu/in-notes/rfc860.txt;   ftp://ftp.isi.
edu/in-notes/std/std31.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc860.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std31.txt. See

also RFC0860 [?].

STD0032

[32] J. Postel and J. Reynolds. STD 32: Telnet Extended Options —

List Option, May 1983. URL ftp://ftp.isi.edu/in-notes/rfc861.
txt;   ftp://ftp.isi.edu/in-notes/std/std32.txt;   ftp://ftp.

math.utah.edu/pub/rfc/rfc861.txt;   ftp://ftp.math.utah.edu/pub/

rfc/std/std32.txt. See also RFC0861 [?].

STD0033

[33] K. Sollins. STD 33: The TFTP Protocol (Revision 2), July 1992.

URL ftp://ftp.isi.edu/in-notes/rfc1350.txt;   ftp://ftp.isi.
edu/in-notes/std/std33.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc1350.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std33.txt.

See also RFC1350 [?].

STD0034

[34] C. Hedrick. STD 34: Routing Information Protocol, June 1988.

URL ftp://ftp.isi.edu/in-notes/rfc1058.txt;   ftp://ftp.isi.
edu/in-notes/std/std34.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc1058.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std34.txt.

See also RFC1058 [?].

STD0035

[35] Marshall T. Rose and Dwight E. Cass. STD 35: ISO Transport Service on top of the TCP (Version: 3), May 1987. URL ftp://ftp.isi.edu/

in-notes/rfc1006.txt;   ftp://ftp.isi.edu/in-notes/std/std35.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1006.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std35.txt. See also RFC1006 [?].

STD0036

[36] D. Katz. STD 36: Transmission of IP and ARP over FDDI net-

works, January 1993. URL ftp://ftp.isi.edu/in-notes/rfc1390.
txt;   ftp://ftp.isi.edu/in-notes/std/std36.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc1390.txt;   ftp://ftp.math.utah.edu/
pub/rfc/std/std36.txt. See also RFC1390 [?].

STD0037

[37] David C. Plummer. STD 37: An Ethernet Address Resolution Protocol — or — Converting Network Protocol Addresses to 48.bit Ethernet Address for Transmission on Ethernet Hardware, November 1982.

URL ftp://ftp.isi.edu/in-notes/rfc826.txt;   ftp://ftp.isi.
edu/in-notes/std/std37.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc826.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std37.txt. See

also RFC0826 [?].

STD0038

[38] Ross Finlayson, Timothy Mann, Jeffrey Mogul, and Marvin Theimer. STD 38: A Reverse Address Resolution Protocol, June 1984. URL ftp://ftp.

isi.edu/in-notes/rfc903.txt;   ftp://ftp.isi.edu/in-notes/std/
std38.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc903.txt;   ftp:
//ftp.math.utah.edu/pub/rfc/std/std38.txt. See also RFC0903 [?].

STD0039

[39] BBN. STD 39: Interface Message Processor: Specifications for the interconnection of a host and an IMP (revised), December 1981. Was BBN Report 1822 (IMP/Host Interface).

STD0040

[40] Bolt Beranek and Newman Laboratories. STD 40: Host Access Protocol specification, July 1984. URL ftp://ftp.isi.edu/in-notes/

rfc907.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc907.txt. See

also RFC0907 [?].

STD0041

[41] Charles Hornig. STD 41: A standard for the transmission of IP datagrams over Ethernet networks, April 1984. URL ftp://ftp.isi.edu/in-

notes/rfc894.txt;   ftp://ftp.isi.edu/in-notes/std/std41.txt;
 ftp://ftp.math.utah.edu/pub/rfc/rfc894.txt;   ftp://ftp.math.
utah.edu/pub/rfc/std/std41.txt. See also RFC0894 [?].

STD0042

[42] Jon Postel. STD 42: A standard for the transmission of IP datagrams over experimental Ethernet networks, April 1984. URL ftp://ftp.isi.edu/

in-notes/rfc895.txt;   ftp://ftp.isi.edu/in-notes/std/std42.txt;
 ftp://ftp.math.utah.edu/pub/rfc/rfc895.txt;   ftp://ftp.math.
utah.edu/pub/rfc/std/std42.txt. See also RFC0895 [?].

STD0043

[43] J. Postel and J. Reynolds. STD 43: A standard for the transmission of IP datagrams over IEEE 802 networks, February 1988.

URL ftp://ftp.isi.edu/in-notes/rfc1042.txt;   ftp://ftp.isi.
edu/in-notes/rfc948.txt;   ftp://ftp.isi.edu/in-notes/std/std43.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1042.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc948.txt;   ftp://ftp.math.utah.edu/pub/

rfc/std/std43.txt. See also RFC1042 [?]. Obsoletes RFC0948 [?].

STD0044

[44] D. L. Mills. STD 44: DCN local-network protocols, August 1993.

URL ftp://ftp.isi.edu/in-notes/rfc891.txt;   ftp://ftp.isi.
edu/in-notes/std/std44.txt;   ftp://ftp.math.utah.edu/pub/rfc/

rfc891.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std44.txt. See

also RFC0891 [?].

STD0045

[45] K. Hardwick and J. Lekashman. STD 45: Internet Protocol on Network Systems’ HYPERchannel: Protocol specification, August 1993.

URL ftp://ftp.isi.edu/in-notes/rfc1044.txt;   ftp://ftp.isi.
edu/in-notes/std/std45.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc1044.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std45.txt.

See also RFC1044 [?].

STD0046

[46] D. Provan. STD 46: Transmitting IP traffic over ARCNET net-

works, August 1993. URL ftp://ftp.isi.edu/in-notes/rfc1051.txt;
 ftp://ftp.isi.edu/in-notes/rfc1201.txt;   ftp://ftp.isi.edu/
in-notes/std/std46.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1051.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1201.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std46.txt. See also RFC1201 [?]. Obso-

letes RFC1051 [?].

STD0047

[47] J. L. Romkey. STD 47: A nonstandard for transmission of IP datagrams over serial lines: SLIP, June 1988. URL ftp://ftp.isi.edu/

in-notes/rfc1055.txt;   ftp://ftp.isi.edu/in-notes/std/std47.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1055.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std47.txt. See also RFC1055 [?].

STD0048

[48] L. McLaughlin III. STD 48: A standard for the transmission of IP datagrams over NetBIOS networks, August 1993. URL ftp://ftp.isi.edu/

in-notes/rfc1088.txt;   ftp://ftp.isi.edu/in-notes/std/std48.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1088.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std48.txt. See also RFC1088 [?].

STD0049

[49] L. McLaughlin III. STD 49: A standard for the transmission of 802.2 packets over IPX networks, August 1993. URL ftp://ftp.isi.edu/

in-notes/rfc1132.txt;   ftp://ftp.isi.edu/in-notes/std/std49.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1132.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std49.txt. See also RFC1132 [?].

STD0050

[50] Anonymous. Reserved for definitions of managed objects for the Ethernetlike interface types, July 1994. URL ftp://ftp.isi.edu/in-notes/

rfc3638.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc3638.txt. See

RFC3638 [?].

STD0051

[51] W. Simpson. STD 51: The Point-to-Point Protocol (PPP), July 1994.

URL ftp://ftp.isi.edu/in-notes/rfc1549.txt;   ftp://ftp.isi.
edu/in-notes/rfc1661.txt;   ftp://ftp.isi.edu/in-notes/rfc1662.
txt;   ftp://ftp.isi.edu/in-notes/std/std51.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc1549.txt;   ftp://ftp.math.utah.edu/
pub/rfc/rfc1661.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1662.
txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std51.txt. See also

RFC1661, RFC1662 [?, ?]. Obsoletes RFC1549 [?].

STD0052

[52] D. Piscitello and J. Lawrence. STD 52: The transmission of IP datagrams over the SMDS service, March 1991. URL ftp://ftp.isi.edu/

in-notes/rfc1209.txt;   ftp://ftp.isi.edu/in-notes/std/std52.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1209.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std52.txt. See also RFC1209 [?].

STD0053

[53] J. Myers and M. Rose. STD 53: Post Office Protocol — Ver-

sion 3, May 1996. URL ftp://ftp.isi.edu/in-notes/rfc1725.txt;
 ftp://ftp.isi.edu/in-notes/rfc1939.txt;   ftp://ftp.isi.edu/
in-notes/std/std53.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1725.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1939.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std53.txt. See also RFC1939 [?]. Obso-

letes RFC1725 [?].

STD0054

[54] J. Moy. STD 54: OSPF version 2, April 1998. URL ftp://ftp.isi.edu/

in-notes/rfc2328.txt;   ftp://ftp.isi.edu/in-notes/std/std54.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc2328.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std54.txt. See also RFC2328 [?].

STD0055

[55] C. Brown and A. Malis. STD 55: Multiprotocol interconnect over frame re-

lay, September 1998. URL ftp://ftp.isi.edu/in-notes/rfc1294.txt;

 ftp://ftp.isi.edu/in-notes/rfc1490.txt;   ftp://ftp.isi.edu/
in-notes/rfc2427.txt;   ftp://ftp.isi.edu/in-notes/std/std55.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1294.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc1490.txt;   ftp://ftp.math.utah.edu/
pub/rfc/rfc2427.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std55.

txt. See also RFC2427 [?]. Obsoletes RFC1490, RFC1294 [?, ?].

STD0056

[56] G. Malkin. STD 56: RIP version 2, November 1998. URL ftp://ftp.isi.

edu/in-notes/rfc1723.txt;   ftp://ftp.isi.edu/in-notes/rfc2453.
txt;   ftp://ftp.isi.edu/in-notes/std/std56.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc1723.txt;   ftp://ftp.math.utah.edu/
pub/rfc/rfc2453.txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std56.

txt. See also RFC2453 [?]. Obsoletes RFC1723 [?].

STD0057

[57] G. Malkin. STD 57: RIP Version 2 Protocol Applicability State-

ment, November 1994. URL ftp://ftp.isi.edu/in-notes/rfc1722.
txt;   ftp://ftp.isi.edu/in-notes/std/std57.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc1722.txt;   ftp://ftp.math.utah.edu/
pub/rfc/std/std57.txt. See also RFC1722 [?].

STD0058

[58] K. McCloghrie, D. Perkins, and J. Schoenwaelder. STD 58: Structure of Management Information version 2 (SMIv2), April 1999.

URL ftp://ftp.isi.edu/in-notes/rfc1902.txt;   ftp://ftp.isi.
edu/in-notes/rfc2578.txt;   ftp://ftp.isi.edu/in-notes/rfc2579.
txt;   ftp://ftp.isi.edu/in-notes/rfc2580.txt;   ftp://ftp.isi.
edu/in-notes/std/std58.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc1902.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc2578.txt;
ftp://ftp.math.utah.edu/pub/rfc/rfc2579.txt;   ftp://ftp.math.
utah.edu/pub/rfc/rfc2580.txt;   ftp://ftp.math.utah.edu/pub/rfc/

std/std58.txt. See also RFC2578, RFC2579, RFC2580 [?, ?, ?]. Obsoletes RFC1902 [?].

STD0059

[59] S. Waldbusser. STD 59: Remote network monitoring management informa-

tion base, May 2000. URL ftp://ftp.isi.edu/in-notes/rfc1757.txt;
 ftp://ftp.isi.edu/in-notes/rfc2819.txt;   ftp://ftp.isi.edu/
in-notes/std/std59.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1757.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc2819.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std59.txt. See also RFC2819 [?]. Obso-

letes RFC1757 [?].

STD0060

[60] N. Freed. STD 60: SMTP service extension for command pipelin-

ing, September 2000. URL ftp://ftp.isi.edu/in-notes/rfc2197.txt;
 ftp://ftp.isi.edu/in-notes/rfc2920.txt;   ftp://ftp.isi.edu/
in-notes/std/std60.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc2197.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc2920.txt;   ftp://ftp.
math.utah.edu/pub/rfc/std/std60.txt. See also RFC2920 [?]. Obso-

letes RFC2197 [?].

STD0061

[61] N. Haller, C. Metz, P. Nesser, and M. Straw. STD 61: A one-time password system, February 1998. URL ftp://ftp.isi.edu/in-notes/

rfc1938.txt;   ftp://ftp.isi.edu/in-notes/rfc2289.txt;   ftp:
//ftp.isi.edu/in-notes/std/std61.txt;   ftp://ftp.math.utah.edu/
pub/rfc/rfc1938.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc2289.
txt;   ftp://ftp.math.utah.edu/pub/rfc/std/std61.txt. See also

RFC2289 [?]. Obsoletes RFC1938 [?].

STD0062

[62] D. Harrington, R. Presuhn, and B. Wijnen. STD 62: Simple network management protocol version 3 (SNMPv3), December 2002.

URL ftp://ftp.isi.edu/in-notes/rfc1905.txt;   ftp://ftp.isi.
edu/in-notes/rfc1906.txt;   ftp://ftp.isi.edu/in-notes/rfc1907.
txt;   ftp://ftp.isi.edu/in-notes/rfc2571.txt;   ftp://ftp.isi.
edu/in-notes/rfc2572.txt;   ftp://ftp.isi.edu/in-notes/rfc2573.
txt;   ftp://ftp.isi.edu/in-notes/rfc2574.txt;   ftp://ftp.isi.
edu/in-notes/rfc2575.txt;   ftp://ftp.isi.edu/in-notes/rfc3411.
txt;   ftp://ftp.isi.edu/in-notes/rfc3412.txt;   ftp://ftp.isi.
edu/in-notes/rfc3413.txt;   ftp://ftp.isi.edu/in-notes/rfc3414.
txt;   ftp://ftp.isi.edu/in-notes/rfc3415.txt;   ftp://ftp.isi.
edu/in-notes/rfc3416.txt;   ftp://ftp.isi.edu/in-notes/rfc3417.
txt;   ftp://ftp.isi.edu/in-notes/rfc3418.txt;   ftp://ftp.isi.
edu/in-notes/std/std62.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc1905.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc1906.txt;
ftp://ftp.math.utah.edu/pub/rfc/rfc1907.txt;   ftp://ftp.math.
utah.edu/pub/rfc/rfc2571.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc2572.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc2573.txt;
ftp://ftp.math.utah.edu/pub/rfc/rfc2574.txt;   ftp://ftp.math.
utah.edu/pub/rfc/rfc2575.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc3411.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc3412.txt;
ftp://ftp.math.utah.edu/pub/rfc/rfc3413.txt;   ftp://ftp.math.
utah.edu/pub/rfc/rfc3414.txt;   ftp://ftp.math.utah.edu/pub/rfc/
rfc3415.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc3416.txt;

ftp://ftp.math.utah.edu/pub/rfc/rfc3417.txt;   ftp://ftp.math.
utah.edu/pub/rfc/rfc3418.txt;   ftp://ftp.math.utah.edu/pub/rfc/

std/std62.txt. See also RFC3411, RFC3412, RFC3413, RFC3414, RFC3415, RFC3416, RFC3417, RFC3418 [?, ?, ?, ?, ?, ?, ?, ?]. Obsoletes RFC2571, RFC2572, RFC2573, RFC2574, RFC2575, RFC1905, RFC1906, RFC1907 [?, ?, ?, ?, ?, ?, ?, ?].

STD0063

[63] F. Yergeau. STD 63: UTF-8, a transformation format of ISO 10646,

November 2003. URL ftp://ftp.isi.edu/in-notes/std/std63.txt;
 ftp://ftp.isi.edu/in-notes/rfc2279.txt;   ftp://ftp.isi.edu/
in-notes/rfc3629.txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc2279.
txt;   ftp://ftp.math.utah.edu/pub/rfc/rfc3629.txt. See also

RFC3629 [?]. Obsoletes RFC2279 [?].

STD0064

[64] H. Schulzrinne, S. Casner, R. Frederick, and V. Jacobson. STD 64: RTP: A transport protocol for real-time applications, July 2003.

URL ftp://ftp.isi.edu/in-notes/rfc1889.txt,   ftp://ftp.isi.
edu/in-notes/rfc3550.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1889.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc3550.txt. See also

RFC3550 [?]. Obsoletes RFC1889 [?].

STD0065

[65] H. Schulzrinne and S. Casner. STD 65: RTP profile for audio and video conferences with minimal control, July 2003. URL ftp://ftp.isi.edu/in-

notes/rfc1890.txt,   ftp://ftp.isi.edu/in-notes/rfc3551.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1890.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc3551.txt. See also RFC3551 [?]. Obsoletes

RFC1890 [?].

STD0066

[66] T. Berners-Lee, R. Fielding, and L. Masinter. STD 66: Uniform Resource Identifier (URI): Generic syntax, January 2005. URL ftp://ftp.isi.edu/

in-notes/rfc1738.txt,   ftp://ftp.isi.edu/in-notes/rfc1808.txt,
 ftp://ftp.isi.edu/in-notes/rfc2396.txt,   ftp://ftp.isi.edu/
in-notes/rfc2732.txt,   ftp://ftp.isi.edu/in-notes/rfc3986.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1738.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1808.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2396.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2732.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc3986.txt. See also RFC3986 [?].

Obsoletes RFC2732, RFC2396, RFC1808 [?, ?, ?]. Updates RFC1738 [?].